Circular No. 8753

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 S4 (CHRISTENSEN) E. J. Christensen reports the discovery of another comet on Catalina

Sky Survey images obtained with the 0.68-m Schmidt reflector (discovery observation tabulated below), the object displaying a 1′.8 fan-shaped tail spanning p.a. ≈ 210^{o}–250^{o} in a stacked image of four 30-s exposures obtained in poor seeing. Following posting on the ‘NEO Confirmation Page’, M. Tichý and J. Tichá write that their CCD images taken on Sept. 22.99 UT with the 1.06-m KLENOT Telescope at Kleť show a diffuse 10′′ coma and a tail in p.a. 260^{o}. P. Birtwhistle (Great Shefford, Berkshire, U.K., 0.40-m f/6 Schmidt-Cassegrain reflector) reports that his CCD images taken on Sept. 23.1 show a coma of diameter 10′′ with central condensation and a tail 40′′ long in p.a. 245^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Sept.22.37326  1^{h}50^{m}43.^{s}48  −16^{o}25′58′′.4  16.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-S53.

T = 2007 Apr. 26.414 TT  ω = 28.^{o}144
Ω = 28.811 }2000.0
q = 2.89250 AU  i = 52.670

RS OPHIUCHI M. Orio and T. Nelson, Istituto Nazionale di Astrofisica and University

of Wisconsin, report that RS Oph (cf. IAUC 8671, 8695) was observed with XMM-Newton on Sept. 6 for about 15 hr. The count rate in the range 0.1– 10 keV had decreased by more than a factor of 1000 with respect to the last XMM-Newton observation on Apr. 6. The unabsorbed flux is at least 4 × 10^{−12} erg cm^{−2} s^{−1}. There is still significant continuum flux below 2 keV, but the spectrum shows mainly strong emission lines. With the RGS spectrograph, several H-like and He-like emission lines of Mg, Ne, and Fe are detected (most notably Ne VII at 2.478 nm, O VIII at 1.896 nm, Ne X at 1.213 nm, and Fe XVII at 1.501 nm). The RGS lines are red-shifted with a velocity not exceeding 500 km/s. Despite the lack of significant flux above 2 keV, two emission lines of the iron triplet K(α) complex are detected with EPIC around 6.4–6.7 keV.

2006 September 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green